OpenStudy (anonymous):

Sides of three square rooms measure 13 feet each, and sides of two square rooms measure 15 feet each. Which expression shows the total area of these five rooms? do i multiply 13 and 15 ?

OpenStudy (smarty):

Nope, you take the area of each square separately... So for the first 3 squares with sides of 13 : 13*13=165 165*3 = 495 15*15=225 225*2 = 450 495+450=945